# InstructPLM-mu: 1-Hour Fine-Tuning of ESM2 Beats ESM3 in Protein Mutation Predictions

### v2

@@ 1 Introduction
Deep mutational scanning (DMS) is an experimental technique that systematically measures the functional impact of a vast number of sequence variants for a given protein (Fowler et al., 2014; Fowler & Fields, 2014; Hanning et al., 2022). By introducing and testing millions of mutations, DMS generates high-resolution maps that link sequence changes to functional outcomes. These datasets have become invaluable for understanding sequence-function relationships, guiding protein engineering, and benchmarking computational prediction models. However, due to the high cost and limited throughput of DMS experiments, it is infeasible to apply them broadly across all proteins or mutation types (Fowler & Fields, 2014). To address this limitation, numerous computational methods have been proposed to predict mutational effects based on sequence features (Lin et al., 2023; Marquet et al., 2024), structural features (Su et al., 2023; Zhang et al., 2024; Sun et al., 2025), and evolutionary information (Meier et al., 2021; Weitzman et al., 2025; Tan et al., 2025a; Sun et al., 2024). Among these, multimodal protein language models (PLMs) have demonstrated strong generalization capabilities, leveraging large-scale unlabeled protein sequences to capture evolutionary and biochemical constraints without explicit supervision (Notin et al., 2023).
@@ 2.2 Multimodal Alignment
Multimodal alignment aligns heterogeneous features across modalities to enhance cross-modal understanding and specific task performance (Baltrušaitis et al., 2018). This field has gained prominence with advances in large language and vision models (Alayrac et al., 2022). Key challenges in multimodal learning include Feature Fusion and Training Paradigm, which are critical for model performance (Tong et al., 2024; Li & Tang, 2024).
@@ 4.3 Fine-tuning strategies
Fine-tune Strategy
(Tunable Parameters)
Token-wise Concat Channel-wise Concat
@@ A.1 Training Details
Multi-model Projector. To integrate sequence and structural representations, we employ a lightweight multi-modal projector implemented as a multi-layer perceptron (MLP) with GELU (Hendrycks & Gimpel, 2016) activation. The projector maps the concatenated embeddings from the protein language backbone and the structure encoder into a unified latent space with a dimension of the backbone model. Layer normalization is applied after each hidden layer.

### v3

@@ 1 Introduction
Deep mutational scanning (DMS) is an experimental technique that systematically measures the functional impact of a vast number of sequence variants for a given protein (Fowler et al., 2014; Fowler and Fields, 2014; Hanning et al., 2022). By introducing and testing millions of mutations, DMS generates high-resolution maps that link sequence changes to functional outcomes. These datasets have become invaluable for understanding sequence-function relationships, guiding protein engineering, and benchmarking computational prediction models. However, due to the high cost and limited throughput of DMS experiments, it is infeasible to apply them broadly across all proteins or mutation types (Fowler and Fields, 2014). To address this limitation, numerous computational methods have been proposed to predict mutational effects based on sequence features (Lin et al., 2023; Marquet et al., 2024), structural features (Su et al., 2023; Zhang et al., 2024; Sun et al., 2025), and evolutionary information (Meier et al., 2021; Weitzman et al., 2025; Tan et al., 2025a; Sun et al., 2024). Among these, multimodal protein language models (PLMs) have demonstrated strong generalization capabilities, leveraging large-scale unlabeled protein sequences to capture evolutionary and biochemical constraints without explicit supervision (Notin et al., 2023).
@@ 2.2 Multimodal Alignment
Multimodal alignment aligns heterogeneous features across modalities to enhance cross-modal understanding and specific task performance (Baltrušaitis et al., 2018). This field has gained prominence with advances in large language and vision models (Alayrac et al., 2022). Key challenges in multimodal learning include Feature Fusion and Training Paradigm, which are critical for model performance (Tong et al., 2024; Li and Tang, 2024).
@@ 4.3 Fine-tuning strategies
Fine-tune Strategy (Tunable Parameters) Token-wise Concat Channel-wise Concat
@@ A.1 Training Details
Multi-model Projector. To integrate sequence and structural representations, we employ a lightweight multi-modal projector implemented as a multi-layer perceptron (MLP) with GELU (Hendrycks and Gimpel, 2016) activation. The projector maps the concatenated embeddings from the protein language backbone and the structure encoder into a unified latent space with a dimension of the backbone model. Layer normalization is applied after each hidden layer.